# Style Ambiguity Loss Without Labeled Datasetss

### v2

@@ Abstract
In this work, we explore using the style ambiguity training objective, originally used to approximate creativity, on a diffusion model. However, this objective requires the use of a pretrained classifier and a labeled dataset. We introduce new forms of style ambiguity loss that do not require training a classifier or a labeled dataset, and show that our new methods score higher both on automated metrics and user studies to analyze novelty and appreciation. Code available at https://github.com/jamesBaker361/clipcreate
1 Introduction
Much of generative imagery is trained to mimic the underlying dataset. However, there is also demand for machine-generated images that are not just realistic but artistic, as exemplified by the picture that sold for nearly half a million dollars (Christie’s, 2018) at auction, or the picture used in the sitcom Silicon Valley (AICAN, 2024). While the definition of art is somewhat philosophical and beyond the scope of this paper, it most certainly has to be creative. Creative assets are usually defined as being "novel and useful" (Diedrich et al., 2015). A string of random characters is novel in that we cannot predict the next characters. However, if the purpose of characters is to compose words that compose sentences that communicate a message, then a string of random characters is not useful. Most people would not consider a string of random characters to be creative or art. If the utility of an image is to depict objects and scenes that humans understand and can recognize, then generative models perform well in that regard. Novelty, on the other hand has been underexplored. A breakthrough was the invention of the Creative Adversarial Network (Elgammal et al., 2017), which used a style ambiguity loss to train a generator network to generate images that could not be classified as belonging to a particular style. However, this style ambiguity loss requires a pretrained classifier: Every set of styles or concepts requires training a classifier before even training a model to generate images, which itself takes time and requires a labeled dataset, the collection and curation of which can be expensive. To circumvent these issues, we propose using a classifier for style ambiguity that does not require any additional training and can be easily applied to any dataset, labeled or unlabeled. We demonstrate this by using reinforcement learning to finetune diffusion models with the various style ambiguity losses. Our contributions are as follows:
  • We developed versatile CLIP-based and K-Means-based creative style ambiguity losses that do not require training a separate style classifier on a labeled dataset.
@@ 1 Introduction
2 Related Work
2.1 Creativity
Creative work has been formulated as work having novelty, in that it differs from other similar objects, and also utility, in that it still performs a function (Cropley, 2006). For example, a Corinthian column has elaborate, interesting, unexpected adornments (novelty) but still holds up a building (utility). A distinction can also be made between "P-creativity", where the work is novel to the creator, and "H-creativity" where the work is novel to everyone (Boden, 1990), the latter being of interest. There have been many attempts to make models that can be "H-creative". Such techniques include using genetic algorithms (DiPaola & Gabora, 2008), reconstructing artifacts from novel collections of attributes (Iqbal et al., 2016), and most relevantly to this work, using Generative Adversarial Networks (Elgammal et al., 2017; Nobari et al., 2021).
2.2 Computational Art
One of the first algorithmic approaches dates back to the 1970s with the now primitive AARON (McCorduck, 1991), which was initially only capable of drawing black and white sketches. Generative Adversarial Networks (Goodfellow et al., 2014), or GANs, were some of the first models to be able to create complex, photorealistic images and seemed to have potential to be able to make art. Despite many problems with GANs, such as mode collapse and unstable training (Saxena & Cao, 2023), GANs and further improvements (Arjovsky et al., 2017; Karras et al., 2019; 2018) were state of the art until the introduction of diffusion Sohl-Dickstein et al. (2015). These models, such as IMAGEN (Saharia et al., 2022), DALLE-3 (Betker et al., ) and the Stable Diffusion family (Rombach et al., 2022b; Esser et al., 2024; Podell et al., 2023) have attained widespread adoption due to their ability to generate high-quality images from text prompts.
2.3 Reinforcement Learning
Reinforcement learning (RL) is a method of training a model by having it take actions that generate a reward signal and change the environment, thus changing the impact and availability of future actions Qiang & Zhongli (2011). RL has been used for tasks as diverse as playing board games (Silver et al., 2017), protein design (Lutz et al., 2023), self-driving vehicles (Kiran et al., 2021) and quantitative finance (Sahu et al., 2023). Policy-gradient RL (Sutton et al., 1999) optimizes a policy $\pi$ that chooses which action to take at any given timestep, as opposed to value-based methods that may use a heuristic to determine the optimal choice. Examples of policy gradient methods include Soft Actor Critic (Haarnoja et al., 2018), Deep Deterministic Policy Gradient (Lillicrap et al., 2019) and Trust Region Policy Optimization (Schulman et al., 2017a).
3 Background
3.1 Creative Adversarial Network
A Generative Adversarial Network, or GAN (Goodfellow et al., 2014), consists of two models, a generator and a discriminator. The generator generates samples from noise, and the discriminator detects if the samples are drawn from the real data or generated. During training, the generator is trained to trick the discriminator into classifying generated images as real, and the discriminator is trained to classify images correctly. Given a generator $G:\mathbb{R}^{noise}\rightarrow\mathbb{R}^{h\times w\times 3}$, a discriminator $D:\mathbb{R}^{h\times w\times 3}\rightarrow[0,1]$ real images $x\in\mathbb{R}^{h\times w\times 3}$, and noise $\mathcal{Z}\in\mathbb{R}^{noise}$, the objective is:
[…]
Elgammal et al. (2017) introduced the Creative Adversarial Network, or CAN, which was a DCGAN (Radford et al., 2016) where the discriminator was also trained to classify real samples, minimizing the style classification loss. Given $N$ classes of image (such as ukiyo-e, baroque, impressionism, etc.), the classification modules of the Discriminator $D_{C}:\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{R}^{N}$ that returns a probability distribution over the $N_{s}$ style classes for an image and the real labels $\ell\in\mathbb{R}^{N}$, the style classification loss was:
@@ 3.2 Diffusion
Once the model has been trained, the reverse process, aka inference, generates a sample from noise $x_{T}\sim\mathcal{N}(0,1)$. We use the DDIM technique (Song et al., 2022) for sampling. In this work, we the more popular variant, Stable Diffusion (Rombach et al., 2022a), where $x$ is replaced with a latent embedding $\mathcal{E}(x)$, where $\mathcal{E}:\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{R}^{h_{z}\times w_{z}\times c_{z}}$ is a frozen autoencoder (Kingma & Welling, 2022), and $h_{z}<h,w_{z}<w,c_{z}>3$.
4 Methods
4.1 Denoising Diffusion Proximal Optimisation
Introduced by Black et al. (2023), Denoising Diffusion Proximal Optimisation, or DDPO, represents the reverse Diffusion Process as a Markov Decision Process (Bellman, 1957). A similar method was also pursued by Fan et al. (2023). Reinforcement learning training was then applied to a pretrained diffusion model. Following Schulman et al. (2017b), Black et al. (2023) also implemented clipping to protect the policy gradient from excessively large updates, and per prompt stat tracking to normalize rewards. We largely follow their method but use a different reward function. We fine-tune off of the pre-existing stabilityai/stable-diffusion-2-base checkpoint (Rombach et al., 2022a) downloaded from https://huggingface.co/stabilityai/stable-diffusion-2-base.
4.2 Text Prompts
@@
4.3 Datasets and Labels
Training the CAN, of course, requires a labeled dataset. We used two different real datasets based off of WikiArt (Saleh & Elgammal, 2015):
@@ 4.3 Datasets and Labels
    Mediums: We used BLIP (Li et al., 2022) to generate captions, and then selected the 10 classes that had the highest fraction of their descriptions containing any of the words in the text prompt set (painting, drawing, art). We then used the images in WikiArt that belonged to those classes. Consists of roughly 20k images.
@@
This also meant we had two different sets of style class labels: $L_{full}$, all 27 class labels, used with Full and $L_{med}$, the 10 labels used in Mediums. For implementation details regarding the captions, and a list of the class labels used for Mediums refer to appendix D
4.4 DDPO Reward Function
@@
In the original DDPO paper, the authors used four different reward functions for four different tasks. For example, they used a scorer trained on the LAION dataset (Schuhmann & Beaumont, 2022) as the reward function to improve the aesthetic quality of generated outputs. In this paper, we use the reward model based on Elgammal et al. (2017), where the model is rewarded for stylistic ambiguity, combined with a reward for utility. Given a pretrained CLIP (Radford et al., 2021) model, that can return a similarity score for each image-text pair: $CLIP:\mathbb{R}^{text}\times\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{R}$, image $x_{0}\in\mathbb{R}^{h\times w\times 3}$ generated with text prompt $s$, cross entropy CE, uniform distribution $U\in\mathbb{R}^{N}$ and a classifier $C:\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{R}^{N}$ we want to maximize:
$R(x_{0})=-\lambda_{novelty}\textbf{CE}(C(x_{0}),U)+\lambda_{utility}CLIP(s,x_{0})$
@@
The first term on the left side of the equation represents style ambiguity loss, and the second term maintains alignment between text and image, essentially keeping the model from straying "too far" from the text prompt; these terms approximate novelty and utility, respectively. We actually have multiple choices of classifier, which we discuss. The Discriminator Classifier is what was used for the CAN, and the CLIP-Based Style Classifier and K-Means Image Based Classifer have never been used for style ambiguity loss before, and our contribution is to introduce them and show their merits.
4.4.1 Discriminator Classifier
@@
We can use the classification module of the CAN discriminator as the classifier in the reward function, setting $C=D_{C}$. This is the traditional method of style ambiguity loss, and the baseline against which we are comparing the other two types of classifier with. This discriminator was trained at image resolution 512, with batch size 128 without using gradient penalty (Gulrajani et al., 2017). Given that we had two datasets (Full and Mediums), we actually had to train two of these discriminators, one trained for each. The disadvantage of using a discriminator is that it requires a labeled dataset, which are expensive and time-consuming to create. Even the use of VLMs (Ghosh et al., 2024) for labeling may not suffice, given that expert domain knowledge is required in many cases. We thus introduce two methods of style ambiguity that do not require a labeled dataset.
4.4.2 CLIP-Based Style Classifier
@@
For each generated image $x_{0}$, for each class name $s_{i},1\leq i\leq N_{s}$ in the style class label set we want to use, we find $CLIP(s_{i},x_{0})$. We can then create a vector $(CLIP(s_{1},x_{0}),CLIP(s_{2},x_{0}),,,CLIP(s_{N_{s}},x_{0}))$ and then use softmax to normalize the vector and define the result as $CC(x_{0})$. Formally:
$CC(x_{0})=\textbf{softmax}((CLIP(s_{1},x_{0}),CLIP(s_{2},x_{0}),,,CLIP(s_{N_{s}},x_{0}))$
@@
Then we set $C=CC$. Given the two sets of labels, $L_{full}$ and $L_{med}$, we had two different types of CLIP-Based classifier.
4.4.3 K-Means Image Based Classifiers
Alternatively, we can embed the images (In our case, we embedded them into the CLIP embedding space $\in\mathbb{R}^{768}$) and perform k-means clustering to generate k centers. Given a CLIP Embedder $E:\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{R}^{768}$ mapping images to embeddings, and the k centers $c_{1},c_{2},,,c_{k}$ we can create a vector $(\frac{1}{||E(x_{0})-c_{1}||},\frac{1}{||E(x_{0})-c_{2}||},,,,\frac{1}{||E(x_{0})-c_{k}||}$ and then use softmax to normalize the vector and define the result as $KM(x_{0}))$. Formally:
$KMEANS(x_{0})=\textbf{softmax}(\frac{1}{||E(x_{0})-c_{1}||},\frac{1}{||E(x_{0})-c_{2}||},,,,\frac{1}{||E(x_{0})-c_{k}||})$
@@ 4.4.3 K-Means Image Based Classifiers
5 Results
Given three types of classifier and two datasets, we had six models.
We generated all images with width and height = 512. The authors used width and height = 256 in the original CAN paper. However, given that larger, more detailed images are preferred by most people, we thought it more relevant to focus on larger images.
5.1 Automated Evaluation
For each dataset, for each choice of classifier, we made an evaluation dataset of 200 images. We used the exact same prompts and random seeds for each. I.e. if the nth image generated by disc-full used prompt "painting" and random seed = $z\in\mathbb{Z}$, so would the nth image generated by clip-full, kmeans-full, etc. for prompts, we had 24 prompts, each generated by concatenating one of ["painting","art","drawing"] with one of [" of a man ", " of a woman "," of nature "," of an animal ", " of a city"," "," of a building "," of a child "] We used two scoring metrics to score the models. Both of these metrics are metrics of human appreciation:
  • AVA Score: (AVA) The AVA model was trained on the AVA dataset (Murray et al., 2016) of images and average rankings by human subjects, in order to learn to approximate human preferences given an image. We used the CLIP model weights from the clip-vit-large-patch14 checkpoint and the Multi-Layer Perceptron weights downloaded from https://huggingface.co/trl-lib/ddpo-aesthetic-predictor.
  • Image Reward: (IR) The image reward model (Xu et al., 2023) was trained to score images given their text description based on a dataset of images and human rankings. We used the image-reward python library found at https://github.com/THUDM/ImageReward/tree/main.
@@ 5.1 Automated Evaluation
$disc-full$ 4.96 (0.5) -1.26 (1.04)
$clip-full$ 4.99 (0.46) -1.59 (0.63)
$kmeans-full$ 5.71 (0.42) 0.03 (1.35)
$disc-med$ 5.12 (0.4) -1.55 (0.59)
$clip-med$ 5.44 (0.41) 0.91 (0.89)
$kmeans-med$ 5.09 (0.67) 0.77 (0.87)
Table 1: Aesthetic Scores, Diffusion Models
[Uncaptioned image] [Uncaptioned image]
Table 2: Score Box Plots
Results of our experiments are shown in Table 1 Each cell contains the mean (and standard deviation). Box plots are shown in 2. We can see that there was less variance and generally less skew for the aesthetic score than for the Image Reward. For both datasets, for both metrics, either a clip model or kmeans model scored higher than the baseline discriminator model. This shows that our discriminator-free approaches in do not come at the cost of lower quality, and in fact are likely to produce better images than the discriminator baseline.
5.1.1 Comparison with GAN Baseline
GANs are largely out of fashion, but we nonetheless thought it worthwhile to test a few variations of the Creative Adversarial Network, as described in Elgammal et al. (2017). Lacking an official repo or weights, we had to base our implementation off of what was described in the paper. The diffusion models disc-full and disc-med used the classification heads of discriminators trained in this fashion for their reward model, but the images being evaluated in this section were created by the generators used in the Creative Adversarial Network. In the original CAN paper, all images were generated at resolution 256, trained on the Full wikiart dataset, with batch size 128 without gradient penalty. However, we wanted to slightly expand our search space, so we tried every possible combination of four choices

### v3

# Style Ambiguity Loss Using CLIP

@@ Abstract
In this work, we explore using the style ambiguity training objective, originally used to approximate creativity, on a diffusion model. However, this objective requires the use of a pretrained classifier and a labeled dataset. We introduce new forms of style ambiguity loss that do not require training a new classifier or a labeled dataset. Instead of using a classifier, we generate centroids in the CLIP embedding space, and images are classified based on their relative distance to said centroids. We find the centroids via K-means clustering of an unlabeled dataset, as well as using text labels to generate CLIP embeddings, to be used as centroids. Code is available at https://github.com/jamesBaker361/clipcreate
Index Terms:
CLIP, multimodal, diffusion
I Introduction
Much of generative imagery is trained to mimic the underlying dataset. Generating truly ”creative” work was initiated by the Creative Adversarial Network [1], which used a style ambiguity loss to train a generator network to generate images that could not be classified as belonging to a particular style. However, every set of styles or concepts requires training a classifier before even training a model to generate images, which itself takes time and requires a labeled dataset, the collection and curation of which can be expensive. To circumvent these issues, we propose using a classification technique for style ambiguity that does not require any additional training and can be easily applied to any dataset, labeled or unlabeled. We demonstrate this by using reinforcement learning to finetune diffusion models with the various style ambiguity losses. Our contributions are as follows:
  • We developed versatile Text-based and K-Means-based creative style ambiguity losses that do not require training a separate style classifier on a labeled dataset, instead using CLIP for generating centroids.
@@ I Introduction
II Background
II-A Creative Adversarial Network
A Generative Adversarial Network, or GAN [2], consists of two models, a generator and a discriminator. The generator generates samples from noise, and the discriminator detects if the samples are drawn from the real data or generated. During training, the generator is trained to trick the discriminator into classifying generated images as real, and the discriminator is trained to classify images correctly. Given a generator $G:\mathbb{R}^{noise}\mathbb{R}ightarrow\mathbb{R}^{h\times w\times 3}$, a discriminator $D:\mathbb{R}^{h\times w\times 3}\mathbb{R}ightarrow[0,1]$ real images $x\in\mathbb{R}^{h\times w\times 3}$, and noise $\mathcal{Z}\in\mathbb{R}^{noise}$, the objective is:
$\underset{G}{min}\>\underset{D}{max}\>\mathbb{E}_{x}[log(D(x)]+\mathbb{E}_{\mathcal{Z}}[log(1-D(G(\mathcal{Z}))]$
[1] introduced the Creative Adversarial Network, or CAN, which was a DCGAN [3] where the discriminator was also trained to classify real samples, minimizing the style classification loss. Given $N$ classes of image (such as ukiyo-e, baroque, impressionism, etc.), the classification modules of the Discriminator $D_{C}:\mathbb{R}^{h\times w\times 3}\mathbb{R}ightarrow\mathbb{R}^{N}$ that returns a probability distribution over the $N_{s}$ style classes for an image and the real labels $\ell\in\mathbb{R}^{N}$, the style classification loss was:
@@ II-B Diffusion
Once the model has been trained, the reverse process, aka inference, generates a sample from noise $x_{T}\sim\mathcal{N}(0,1)$. We use the DDIM technique [4] for sampling. In this work, we the more popular variant, Stable Diffusion [5], where $x$ is replaced with a latent embedding $\mathcal{E}(x)$, where $\mathcal{E}:\mathbb{R}^{h\times w\times 3}\mathbb{R}ightarrow\mathbb{R}^{h_{z}\times w_{z}\times c_{z}}$ is a frozen autoencoder [6], and $h_{z}<h,w_{z}<w,c_{z}>3$.
III Methods
III-A Denoising Diffusion Proximal Optimisation
Introduced by [7], Denoising Diffusion Proximal Optimisation, or DDPO, represents the reverse Diffusion Process as a Markov Decision Process [8]. A similar method was also pursued by [9]. Reinforcement learning training was then applied to a pretrained diffusion model. Following [10], [7] also implemented clipping to protect the policy gradient from excessively large updates, and per prompt stat tracking to normalize rewards. We largely follow their method but use a different reward function. We fine-tune off of the pre-existing stabilityai/stable-diffusion-2-base checkpoint [5] downloaded from https://huggingface.co/stabilityai/stable-diffusion-2-base.
III-B Text Prompts
@@
III-C Datasets and Labels
Training the CAN, of course, requires a labeled dataset. We used two different real datasets based off of WikiArt [11]:
@@ III-C Datasets and Labels
    Mediums: We used BLIP [12] to generate captions, and then selected the 10 classes that had the highest fraction of their descriptions containing any of the words in the text prompt set (painting, drawing, art). We then used the images in WikiArt that belonged to those classes. Consists of roughly 20k images.
@@
This also meant we had two different sets of style class labels: $L_{full}$, all 27 class labels, used with Full and $L_{med}$, the 10 labels used in Mediums.
III-D DDPO Reward Function
@@
In the original DDPO paper, the authors used four different reward functions for four different tasks. In this paper, we use the reward model based on stylistic ambiguity, combined with a reward for utility. Given a pretrained CLIP [13] model, that can return a similarity score for each image-text pair: $CLIP:\mathbb{R}^{text}\times\mathbb{R}^{h\times w\times 3}\mathbb{R}ightarrow\mathbb{R}$, image $x_{0}\in\mathbb{R}^{h\times w\times 3}$ generated with text prompt $s$, cross entropy CE, uniform distribution $U\in\mathbb{R}^{N}$ and a classifier $C:\mathbb{R}^{h\times w\times 3}\mathbb{R}ightarrow\mathbb{R}^{N}$ we want to maximize:
$R(x_{0})=\lambda_{novelty}\textbf{CE}(C(x_{0}),U)+\lambda_{utility}CLIP(s,x_{0})$
@@
The first term on the left side of the equation represents style ambiguity loss, and the second term maintains alignment between text and image, essentially keeping the model from straying ”too far” from the text prompt; these terms approximate novelty and utility, respectively. We actually have multiple choices of classifier, which we discuss. The Discriminator Classifier is what was used for the CAN, and the Text-Based Style Classifier and K-Means Image Based Classifer have never been used for style ambiguity loss before, and our contribution is to introduce them and show their merits. Many other works [14, 15] have incorporated CLIP embeddings as a condition to generate images, but we only use the CLIP classifier to generate a reward.
III-D1 Discriminator Classifier
@@
We can use the classification module of the CAN discriminator as the classifier in the reward function, setting $C=D_{C}$. This is the baseline against which we are comparing the other two types of classifier with. This discriminator was trained at image resolution 512, with batch size 128. Given that we had two datasets (Full and Mediums), we actually had to train two of these discriminators, one trained for each.
III-D2 Text-Based Style Classifier
@@
For each generated image $x_{0}$, for each class name $s_{i},1\leq i\leq N_{s}$ in the style class label set we want to use, we find $CLIP(s_{i},x_{0})$. We can then create a vector $(CLIP(s_{1},x_{0}),CLIP(s_{2},x_{0}),,,CLIP(s_{N_{s}},x_{0}))$ and then use softmax to normalize the vector and define the result as $TextC(x_{0})$. Formally:
$TextC(x_{0})=\textbf{softmax}((CLIP(s_{1},x_{0}),,,CLIP(s_{N_{s}},x_{0}))$
@@
Then we set $C=TextC$. Given the two sets of labels, $L_{full}$ and $L_{med}$, we had two different types of Text-Based classifier.
III-D3 K-Means Image Based Classifiers
[…]
@@ III-D3 K-Means Image Based Classifiers
IV Results
Given three types of classifier and two datasets, we had six models. Each model was trained for 100 epochs on a single NVIDIA A100 Tensor Core GPU, elapsing roughly 14 hours each. Note that for methods text-full and text-med, we don’t use the data in the data for the classifier per se; however, we use the style class labels, which are unique to the datasets in question. text-full uses $L_{full}$ and text-med uses $L_{mediums}$. We generated all images with width and height = 512.
IV-A Automated Evaluation
For each dataset, for each choice of classifier, we made an evaluation dataset of 200 images. We used the exact same prompts and random seeds for each. I.e. if the nth image generated by disc-full used prompt ”painting” and random seed = $z\in\mathbb{Z}$, so would the nth image generated by text-full, kmeans-full, etc. for prompts, we had 24 prompts, each generated by concatenating one of [”painting”,”art”,”drawing”] with one of [” of a man ”, ” of a woman ”,” of nature ”,” of an animal ”, ” of a city”,” ”,” of a building ”,” of a child ”] We used two scoring metrics to score the models. Both of these metrics are metrics of human appreciation:
  • AVA Score: (AVA) The AVA model was trained on the AVA dataset [16] of images and average rankings by human subjects, in order to learn to approximate human preferences given an image. We used the CLIP model weights from the clip-vit-large-patch14 checkpoint and the Multi-Layer Perceptron weights downloaded from https://huggingface.co/trl-lib/ddpo-aesthetic-predictor.
  • Image Reward: (IR) The image reward model [17] was trained to score images given their text description based on a dataset of images and human rankings. We used the image-reward python library found at https://github.com/THUDM/ImageReward/tree/main.
@@ IV-A Automated Evaluation
discfulldisc-full 4.96 (0.5) -1.26 (1.04)
clipfullclip-full 4.99 (0.46) -1.59 (0.63)
kmeansfullkmeans-full 5.71 (0.42) 0.03 (1.35)
discmeddisc-med 5.12 (0.4) -1.55 (0.59)
clipmedclip-med 5.44 (0.41) 0.91 (0.89)
kmeansmedkmeans-med 5.09 (0.67) 0.77 (0.87)
TABLE I: Aesthetic Scores, Diffusion Models
Results of our experiments are shown in Table I Each cell contains the mean (and standard deviation). For both datasets, for both metrics, either a text model or kmeans model scored higher than the baseline discriminator model. This shows that our discriminator-free approaches in do not come at the cost of lower quality, and in fact are likely to produce better images than the discriminator baseline.
IV-B User Study
We also conducted a user study to evaluate a subset of the models. The goal of this study was to evaluate each model for human appreciation, as we had done with IR and AVA, as well as novelty, surprise, ambiguity and complexity. For each model we did evaluate in this user study, we generated five images with the same seeds. Then, for each image, users to gauge 5 metrics, by asking them each a question, as was done in [1]. These 5 metrics were based on the components of aesthetic arousal [18], that being the level of alertness or excitation that an image provokes in a viewer. ”Too much” arousal can actually have a negative effect on subjective appreciation. The metrics and their associated questions are as follows: